If there were 500 students in Jamal’s class, approximately how many actual students scored higher than Jamal on the quiz if Jamal had a z-score of -1?

  1. 250
  2. 340
  3. 420
  4. 490
Answer #1

Solutions: N= 5oo. P (student score higher than formal) = P/a>-) = 1-P12E-1) = 0.8413 so number of students schored higher th